# Handover Delay Minimization in Non-Terrestrial Networks: Impact of Open RAN Functional Splits

### v1

@@ Abstract
This paper addresses the challenge of optimizing handover (HO) performance in non-terrestrial networks (NTNs) to enhance user equipment (UE) effective service time, defined as the active service time excluding HO delays and radio link failure (RLF) periods. Availability is defined as the normalized effective service time which is effected by different HO scenarios: Intra-satellite HO is the HO from one beam to another within the same satellite; inter-satellite HO refers to the HO from one satellite to another where satellites can be connected to the same or different GSs. We investigate the impact of open radio access network (O-RAN) functional splits (FSs) between ground station (GS) and LEO satellites on HO delay and assess how beam configurations affect RLF rates and intra- and inter-satellite HO rates. This work focuses on three O-RAN FSs—split 7.2x (low layer 1 functions on the satellite), split 2 (layer 1 and layer 2 functions on the satellite), and gNB onboard the satellite—and two beam configurations (19-beam and 127-beam). In a realistic dynamic LEO satellite constellation where different types of HO scenarios are simulated, we maximize effective service time by tuning the time-to-trigger (TTT) and HO margin (HOM) parameters. Our findings reveal that the gNB onboard the satellite achieves the highest availability, approximately $95.4$%, while the split 7.2x exhibits the lowest availability, around $92.8$% due to higher intra-satellite HO delays.
Index Terms:
non-terrestrial network (NTN), conditional handover (CHO), open radio access network (O-RAN), low earth orbit (LEO) satellite, radio link failure (RLF)
@@ II-A System Architecture
As shown in Figure 1, we consider a ground-based UE served by LEO satellites with O-RAN FSs. Each satellite is equipped with highly directional circular aperture antennas [16] that serve ground UEs using Earth-moving beams (non-steerable beams). The visibility of the UE and GS is determined by their respective elevation angles. The elevation angle is defined as the angle between the horizontal plane at the UE or GS and the line connecting it to the satellite. A UE is considered visible to the satellite if its elevation angle exceeds a predefined minimum threshold, $\theta_{min}$, ensuring an unobstructed line-of-sight (LoS) as shown in Figure 1. Similarly, GS visibility is constrained by a minimum GS elevation angle, $\beta_{min}$. Within this overlapped visibility region, we analyze CHO performance for satellites as they move in and out over time. Let the LEO satellite index be denoted as $s\in\{1,2,\ldots,\text{S}\}$, beam indices be denoted as $i\in\{1,2,\ldots,\text{I}\}$ and the GS indices be denoted as $j\in\{1,2,\ldots,\text{J}\}$. Time slots within the service period are indicated by $t\in\{1,2,\ldots,\text{T}\}$ and time slot duration is defined as $\Delta t$. The service time of the UE is defined as $\text{T}\Delta t$. The GS association with satellite $s$ through feeder link (FL) at time $t$ is indicated by $g_{s}(t)$, where $g_{s}(t)$ can take values from set {0,1,2,…,J}. Here, 00 indicates no association of satellite $s$ with any GS. Additionally, at time $t$ the elevation angle between UE and satellite $s$ and between GS $j$ and satellite $s$ is denoted by $\theta_{s}(t)$ and $\beta_{s,j}(t)$ respectively. Furthermore, during each time slot, the UE has access to a set of candidate satellites that meet specific visibility conditions. The candidate satellite set $\mathcal{S}$($t$) at time $t$ is defined as
@@ II-C Conditional Handover Delay Model
where $\tau^{\text{sync}}$ is the synchronization delay, $\tau^{\text{core}}$ is the core network application programming interface (API) call delay, $\tau_{f,q}^{\text{proc}}$ is the processing delay, $\tau_{f,q}^{\text{prop}}(t)$ is the propagation delay, and $\textbf{I}\{\cdot\}$ is an indicator function, equals to $1$ if the condition within parentheses holds and 00 otherwise. The processing delay $\tau_{f,q}^{\text{proc}}$ is modeled as:
$\tau_{f,q}^{\text{proc}}=\ (N_{f,q}^{\text{FL}}+N_{f,q}^{\text{SL}}+N_{f,q}^{\text{ISL}}+N_{f,q}^{\text{IGSL}})\cdot\tau^{\text{ppm}},$ (9)
@@ II-D5 Effective service time
1:  Input: $S$, $I$, $J$, $\Delta t$, $T$, $P_{thr}$, $f$, TTT range, HOM range, $g_{s}(t)$ and $\mathcal{S}(t),\quad\forall t\in\{1,\ldots T\}$,
@@
(a) RSRP $\delta_{t}$=00 s, $\delta_{h}$=00 dB
(b) RSRP $\delta_{t}$=$5$ s, $\delta_{h}$=00 dB
(c) RSRP $\delta_{t}$=00 s, $\delta_{h}$=$3$ dB
(d) Handover $\delta_{t}$=00 s, $\delta_{h}$=00 dB
(e) Handover $\delta_{t}$=$5$ s, $\delta_{h}$=00 dB
(f) Handover $\delta_{t}$=00 s, $\delta_{h}$=$3$ dB
@@ III Problem Formulation
$\tau_{\text{ppm}}$ $1$ ms
UE location ${[}45.78^{\circ}\text{N},1.75^{\circ}\text{E}{]}$
@@
$P_{thr}$ $-120$ dBm
$f_{c}$ $20$ GHz (Ka-Band)
$\delta_{h}$ range [00, $15$] dB
$\delta_{h}$ granularity $5$ dB
$\delta_{t}$ range [00, $30$] s
$\delta_{t}$ granularity $3$ s
@@ IV Results and Analysis
The system-level simulation setup detailed in Section II is implemented following recommendations from [17, 16]. The main parameters for our setup are outlined in Table I. For constellation dynamics, the Starlink LEO configuration is used, comprising $1584$ satellites in $72$ orbital planes, with $22$ satellites per plane at an altitude of $550$ km and inclination of $53^{\circ}$ [19]. Each satellite has beam configurations of $\text{I}=\{19,127\}$, distributed on the ground in $3$ and $6$ concentric tiers as outlined in [10, 11, 16]. Each beam covers a ground area with a diameter of $50$ km and the satellite specifications are based on the assumptions in [17]. Starlink GS locations are incorporated [20], with the number of GSs $\text{J}=4$. The GS positions are: ${[}38.33458^{\circ}\text{N},0.4909^{\circ}\text{E};$ $50.9871^{\circ}\text{N},2.1255^{\circ}\text{E};$ $45.3207^{\circ}\text{N},9.1886^{\circ}\text{E};$ $50.3353^{\circ}\text{N},8.5320^{\circ}\text{E}{]}$. The UE location is positioned to encompass the selected GSs, with coordinates listed in Table I. The satellite associates with a GS upon entering the visibility of the UE, based on the criterion of the highest GS elevation angle. Each GS can simultaneously serve up to two satellites through FLs. If the satellite with the highest elevation angle finds the best GS occupied, it associates with the GS offering the next highest elevation angle. The simulation time is set to $20$ minutes, during which the high density of the LEO constellation ensures a significant number of candidate satellites are visible to the UE at any given moment. Since each satellite remains visible to the UE for around a minute, the $20$-minute duration captures enough satellite interactions to produce reliable results.
@@
(a) RSRP $\delta_{t}$=00 s, $\delta_{h}$=00 dB
(b) RSRP $\delta_{t}$=$5$ s, $\delta_{h}$=00 dB
(c) RSRP $\delta_{t}$=00 s, $\delta_{h}$=$3$ dB
(d) Handover $\delta_{t}$=00 s, $\delta_{h}$=00 dB
(e) Handover $\delta_{t}$=$5$ s, $\delta_{h}$=00 dB
(f) Handover $\delta_{t}$=00 s, $\delta_{h}$=$3$ dB
@@
To further examine the HO behavior in 19 beam configuration, Figure 3(d) illustrates inter-satellite HOs across the simulation, with intra-satellite HOs displayed in an inset plot to provide insight into beam-specific HO patterns. For $\delta_{t}$ = 00 s and $\delta_{h}$ = 00 dB, intra-satellite HOs show a pronounced PP effect, while some inter-satellite HOs result in UHOs. Increasing $\delta_{t}$, in Figure 3(e), reduces the PP effect, and an increased $\delta_{h}$ similarly reduces UHOs, as in Figure 3(f).
@@
In Figure 5(b), intra-satellite HO rates for both configurations are presented under varying HO parameters. At $\delta_{t}$ = 00 s and $\delta_{h}$ = 00 dB, the intra-satellite HO rate is high, approximately $15$ HOs/min for the 19-beam and $10$ HOs/min for the 127-beam, reflecting the frequent PP effect in the former, as shown in Figure 3(d). When $\delta_{t}$ exceeds $5$ s, regardless of $\delta_{h}$, the HO rate in the 127-beam configuration slightly exceeds that of the $19$-beam configuration, as seen in Figure 5(b). This is because the smaller footprint in 19-beam setup may no longer cover the UE, while in the 127-beam setup, the satellite footprint can still maintain coverage, supporting intra-satellite HO even after $5$ s.
@@ Acknowledgement
This work was supported in part by the CELTIC-NEXT Project, 6G for Connected Sky (6G-SKY), with funding received from the Vinnova, Swedish Innovation Agency.

### v2

@@ Abstract
This paper addresses the challenge of optimizing handover (HO) performance in non-terrestrial networks (NTNs) to enhance user equipment (UE) effective service time, defined as the active service time excluding HO delays and radio link failure (RLF) periods. Availability is defined as the normalized effective service time which is effected by different HO scenarios: Intra-satellite HO is the HO from one beam to another within the same satellite; inter-satellite HO refers to the HO from one satellite to another where satellites can be connected to the same or different GSs. We investigate the impact of open radio access network (O-RAN) functional splits (FSs) between ground station (GS) and LEO satellites on HO delay and assess how beam configurations affect RLF rates and intra- and inter-satellite HO rates. This work focuses on three O-RAN FSs—split 7.2x (low layer 1 functions on the satellite), split 2 (layer 1 and layer 2 functions on the satellite), and gNB onboard the satellite—and two beam configurations (19-beam and 127-beam). In a realistic dynamic LEO satellite constellation where different types of HO scenarios are simulated, we maximize effective service time by tuning the time-to-trigger (TTT) and HO margin (HOM) parameters. Our findings reveal that the gNB onboard the satellite achieves the highest availability, approximately 95.495.4%, while the split 7.2x exhibits the lowest availability, around 92.892.8% due to higher intra-satellite HO delays.
@@ II-A System Architecture
As shown in Figure 1, we consider a ground-based UE served by LEO satellites with O-RAN FSs. Each satellite is equipped with highly directional circular aperture antennas [16] that serve ground UEs using Earth-moving beams (non-steerable beams). The visibility of the UE and GS is determined by their respective elevation angles. The elevation angle is defined as the angle between the horizontal plane at the UE or GS and the line connecting it to the satellite. A UE is considered visible to the satellite if its elevation angle exceeds a predefined minimum threshold, $\theta_{min}$, ensuring an unobstructed line-of-sight (LoS) as shown in Figure 1. Similarly, GS visibility is constrained by a minimum GS elevation angle, $\beta_{min}$. Within this overlapped visibility region, we analyze CHO performance for satellites as they move in and out over time. Let the LEO satellite index be denoted as $s\in\{1,2,\ldots,\text{S}\}$, beam indices be denoted as $i\in\{1,2,\ldots,\text{I}\}$ and the GS indices be denoted as $j\in\{1,2,\ldots,\text{J}\}$. Time slots within the service period are indicated by $t\in\{1,2,\ldots,\text{T}\}$ and time slot duration is defined as $\Delta t$. The service time of the UE is defined as $\text{T}\Delta t$. The GS association with satellite $s$ through feeder link (FL) at time $t$ is indicated by $g_{s}(t)$, where $g_{s}(t)$ can take values from set {0,1,2,…,J}. Here, 0 indicates no association of satellite $s$ with any GS. Additionally, at time $t$ the elevation angle between UE and satellite $s$ and between GS $j$ and satellite $s$ is denoted by $\theta_{s}(t)$ and $\beta_{s,j}(t)$ respectively. Furthermore, during each time slot, the UE has access to a set of candidate satellites that meet specific visibility conditions. The candidate satellite set $\mathcal{S}$($t$) at time $t$ is defined as
@@ II-C Conditional Handover Delay Model
where $\tau^{\text{sync}}$ is the synchronization delay, $\tau^{\text{core}}$ is the core network application programming interface (API) call delay, $\tau_{f,q}^{\text{proc}}$ is the processing delay, $\tau_{f,q}^{\text{prop}}(t)$ is the propagation delay, and $\textbf{I}\{\cdot\}$ is an indicator function, equals to $1$ if the condition within parentheses holds and 0 otherwise. The processing delay $\tau_{f,q}^{\text{proc}}$ is modeled as:
$\tau_{f,q}^{\text{proc}}=N_{f,q}^{\text{FL}}\cdot\tau^{\text{ppm}}+N_{f,q}^{\text{SL}}\cdot\tau^{\text{ppm}}+N_{f,q}^{\text{ISL}}\cdot\tau^{\text{ppm}}+N_{f,q}^{\text{IGSL}}\cdot\tau^{\text{ppm}}$ (9)
@@ II-D5 Effective service time
1:Input: $S$, II, $J$, $\Delta t$, $T$, $P_{thr}$, $f$, TTT range, HOM range, $g_{s}(t)$ and $\mathcal{S}(t),\quad\forall t\in\{1,\ldots T\}$,
@@
(a) RSRP $\delta_{t}$=0 s, $\delta_{h}$=0 dB
(b) RSRP $\delta_{t}$=55 s, $\delta_{h}$=0 dB
(c) RSRP $\delta_{t}$=0 s, $\delta_{h}$=33 dB
(d) Handover $\delta_{t}$=0 s, $\delta_{h}$=0 dB
(e) Handover $\delta_{t}$=55 s, $\delta_{h}$=0 dB
(f) Handover $\delta_{t}$=0 s, $\delta_{h}$=33 dB
@@ III Problem Formulation
$\tau_{\text{ppm}}$ 11 ms
UE location $45.78^{\circ}\text{N},1.75^{\circ}\text{E}{]}$
@@
$P_{thr}$ 120-120 dBm
$f_{c}$ 2020 GHz (Ka-Band)
$\delta_{h}$ range [0, $15$] dB
$\delta_{h}$ granularity 55 dB
$\delta_{t}$ range [0, $30$] s
$\delta_{t}$ granularity 33 s
@@ IV Results and Analysis
The system-level simulation setup detailed in Section II is implemented following recommendations from [17, 16]. The main parameters for our setup are outlined in Table I. For constellation dynamics, the Starlink LEO configuration is used, comprising $1584$ satellites in $72$ orbital planes, with $22$ satellites per plane at an altitude of $550$ km and inclination of $53^{\circ}$ [19]. Each satellite has beam configurations of $\text{I}=\{19,127\}$, distributed on the ground in $3$ and $6$ concentric tiers as outlined in [10, 11, 16]. Each beam covers a ground area with a diameter of $50$ km and the satellite specifications are based on the assumptions in [17]. Starlink GS locations are incorporated [20], with the number of GSs $\text{J}=4$. The GS positions are: $38.33458^{\circ}\text{N},0.4909^{\circ}\text{E};$ $50.9871^{\circ}\text{N},2.1255^{\circ}\text{E};$ $45.3207^{\circ}\text{N},9.1886^{\circ}\text{E};$ $50.3353^{\circ}\text{N},8.5320^{\circ}\text{E}{]}$. The UE location is positioned to encompass the selected GSs, with coordinates listed in Table I. The satellite associates with a GS upon entering the visibility of the UE, based on the criterion of the highest GS elevation angle. Each GS can simultaneously serve up to two satellites through FLs. If the satellite with the highest elevation angle finds the best GS occupied, it associates with the GS offering the next highest elevation angle. The simulation time is set to $20$ minutes, during which the high density of the LEO constellation ensures a significant number of candidate satellites are visible to the UE at any given moment. Since each satellite remains visible to the UE for around a minute, the $20$-minute duration captures enough satellite interactions to produce reliable results.
@@
(a) RSRP $\delta_{t}$=0 s, $\delta_{h}$=0 dB
(b) RSRP $\delta_{t}$=55 s, $\delta_{h}$=0 dB
(c) RSRP $\delta_{t}$=0 s, $\delta_{h}$=33 dB
(d) Handover $\delta_{t}$=0 s, $\delta_{h}$=0 dB
(e) Handover $\delta_{t}$=55 s, $\delta_{h}$=0 dB
(f) Handover $\delta_{t}$=0 s, $\delta_{h}$=33 dB
@@
To further examine the HO behavior in 19 beam configuration, Figure 3(d) illustrates inter-satellite HOs across the simulation, with intra-satellite HOs displayed in an inset plot to provide insight into beam-specific HO patterns. For $\delta_{t}$ = 0 s and $\delta_{h}$ = 0 dB, intra-satellite HOs show a pronounced PP effect, while some inter-satellite HOs result in UHOs. Increasing $\delta_{t}$, in Figure 3(e), reduces the PP effect, and an increased $\delta_{h}$ similarly reduces UHOs, as in Figure 3(f).
@@
In Figure 5(b), intra-satellite HO rates for both configurations are presented under varying HO parameters. At $\delta_{t}$ = 0 s and $\delta_{h}$ = 0 dB, the intra-satellite HO rate is high, approximately $15$ HOs/min for the 19-beam and $10$ HOs/min for the 127-beam, reflecting the frequent PP effect in the former, as shown in Figure 3(d). When $\delta_{t}$ exceeds $5$ s, regardless of $\delta_{h}$, the HO rate in the 127-beam configuration slightly exceeds that of the $19$-beam configuration, as seen in Figure 5(b). This is because the smaller footprint in 19-beam setup may no longer cover the UE, while in the 127-beam setup, the satellite footprint can still maintain coverage, supporting intra-satellite HO even after $5$ s.
@@ Acknowledgement
This work was funded by the CELTIC-NEXT Project, 6G for Connected Sky (6G-SKY), with funding received from the Vinnova, Swedish Innovation Agency and was also supported in part by SSF research center, Sustainable Mobile Autonomous and Resilient 6G SatCom (SMART-6GSAT), under registry number CSG23-0001.